### v1

# Bayesian Hierarchical Models for Quantitative Estimates for Performance metrics applied to Saddle Search Algorithms

Abstract
The increasing use of high-throughput computational chemistry demands rigorous methods for evaluating algorithm performance. We present a Bayesian hierarchical modeling paradigm (brms/Stan) for analyzing key performance metrics: function evaluations, computation time, and success/failure. This framework accounts for variability across different systems and functionals, providing reliable uncertainty estimates beyond subjective visual assessments or frequentist limitations. We applied this to compare conjugate gradient (CG) and L-BFGS algorithms for the Dimer method’s rotation phase (EON, with/without removal of external rotations/translations) on a benchmark of 500 initial saddle search approximations, analyzing over 2000 runs. Our results show CG rotations generally outperform L-BFGS, exhibiting a statistically credible, small reduction in PES calls and significantly higher odds of successful convergence. Conversely, enabling rotation removal incurred a substantial PES call penalty without a corresponding credible improvement in success odds in this implementation. These findings, from our novel Bayesian hierarchical modeling application, suggest CG may be preferable for Dimer rotational optimization in similar contexts. This robust statistical framework highlights benefits for revisiting optimization strategies, quantifying uncertainty, and facilitating improved high-throughput computational chemistry methods.
Keywords Bayesian statistics, performance measures
1 Introduction
Locating transition states (TS), first-order saddle points on potential energy surfaces (PES), is fundamental to understanding reaction mechanisms and kinetics in chemistry [10, 35]. Numerous algorithms target this challenge, with minimum mode following (MMF) methods forming a prominent class [7]. These algorithms, initiated from a single point on the PES, iteratively ascend the minimum curvature mode while relaxing orthogonal directions. Notable MMF families include Lanczos-based Activation Relaxation Technique Nouveau (ARTn) methods [29, 5, 30, 36, 13] and the two-point Dimer method [14], which uses a pair of images to estimate the minimum mode direction. Other variations employ internal coordinates [15] or iteratively build PES approximations [21].
Evaluating and comparing such algorithms typically relies on benchmarks, from established sets like Baker’s [2] to more recent collections [8]. However, studies introducing methodological advancements often demonstrate performance on limited sets of small systems, frequently using computationally inexpensive potentials [21, 27, 38, 20, 34, 24]. This traditional benchmarking paradigm appears increasingly misaligned with current computational practice, where peta- and exa-scale resources combined with workflow engines [18, 28] enable large-scale, systematic explorations across diverse chemical spaces. Simple performance comparisons based on small benchmarks or average metrics often neglect significant system-to-system variability and potential implementation biases [22, 11], providing an inadequate basis for selecting optimal methods for demanding high-throughput applications.
To bridge this gap, we advocate for and demonstrate a rigorous statistical framework using Bayesian generalized linear mixed-effects models (GLMMs) for analyzing algorithm performance data from large benchmark sets. This approach, implemented via the brms R package [4] interfacing with Stan [6], robustly handles the statistical modeling of various performance metrics (e.g., computation time, PES call counts, binary success) and explicitly models the hierarchical structure of benchmark data (e.g., multiple runs within chemical systems) using random effects. Crucially, the Bayesian paradigm provides comprehensive uncertainty quantification through full posterior distributions and credible intervals, facilitating nuanced and reliable comparisons beyond simple point estimates or visual inspection.
In this work, we apply this Bayesian GLMM framework to analyze the performance trade-offs associated with two common variations within the Dimer method: the choice between Conjugate Gradient (CG) and Limited-memory BFGS (L-BFGS) optimizers for the rotational step [20], and the effect of enabling or disabling the removal of external rotation and translation degrees of freedom [27]. We utilize the EON software package [9] interfaced with NWChem [1] at the HF/3-21G level of theory, analyzing performance across a diverse benchmark set of 500 initial configurations near saddle points introduced by Hermes et al. [15].
The remainder of this paper is structured as follows: Section 2 details the Dimer method variants and the statistical models employed. Section 3 presents the results of applying these models to the benchmark dataset. Section 4 discusses the implications of the findings and the advantages of the statistical approach, and Section 5 provides concluding remarks.
2 Methods
2.1 Dimer method
The Dimer method [14] is a single-ended saddle search method that algorithmically finds first-order saddle point geometries by iteratively refining the position $R$ (a 3N-dimensional vector representing the system coordinates) and orientation $\hat{N}$ (a unit vector in $3N$ dimensions) of a “dimer”. The dimer consists of two system images, $R_{1}$ and $R_{2}$, separated by a small fixed distance $\Delta R$ along the dimer axis $\hat{N}$:
$R_{1}$ $=R-\frac{\Delta R}{2}\hat{N}$ (1)
$R_{2}$ $=R+\frac{\Delta R}{2}\hat{N}$ (2)
The core idea is to rotate the dimer orientation $\hat{N}$ to align with the lowest curvature mode at the midpoint $R$, and then translate $R$ uphill along this mode (if curvature is negative) while minimizing energy in all other directions. This process typically involves two main steps per iteration: rotation and translation.
The goal of the rotation step is to find the orientation $\hat{N}$ that minimizes the dimer’s energy, keeping the midpoint $R$ fixed. This corresponds to aligning $\hat{N}$ with the direction of lowest local curvature. The rotational dynamics are driven by the component of the sum of forces at the endpoints perpendicular to the dimer axis, $F_{rot}=(F_{1}+F_{2})-((F_{1}+F_{2})\cdot\hat{N})\hat{N}$, where $F_{1}=F(R_{1})$ and $F_{2}=F(R_{2})$. The rotation aims to drive $F_{rot}$ to zero.
Calculating the exact energy curvature requires the Hessian matrix. To avoid this computational expense, the Dimer method typically employs approximations based only on energies and forces at the endpoints $R_{1}$ and $R_{2}$. Following the approach refined by Olsen et al. [34], the curvature along the dimer axis $C(\hat{N})$ and an effective torque or rotational force $\tau(\hat{N})$ can be estimated efficiently and iteratively until a threshold for the rotational force is reached. The finite difference approximation for the curvature is:
$C(\hat{N})\approx\frac{(F_{2}-F_{1})\cdot\hat{N}}{\Delta R}$ (3)
The rotational force $\tau(\hat{N})$, which is perpendicular to $\hat{N}$, can also be estimated from $F_{1}$ and $F_{2}$ [16]. An optimization algorithm, such as Conjugate Gradient (CG) or Limited-memory BFGS (L-BFGS) as investigated in this work, is then used to iteratively update the orientation $\hat{N}$ (typically by small rotations) to minimize the energy with respect to rotation, effectively driving $\tau(\hat{N})$ towards zero.
Once the dimer is aligned with the estimated lowest mode direction $\hat{N}$, the midpoint $R$ is moved towards the saddle point. The translation is guided by an effective force $F_{trans}$. If the estimated curvature $C(\hat{N})$ is negative (indicating an uphill direction towards the saddle), the force component parallel to $\hat{N}$ is inverted. The effective force for the optimizer is:
$F_{trans}(R)=F(R)-2(F(R)\cdot\hat{N})\hat{N}\quad\text{if }C(\hat{N})<0$ (4)
If $C(\hat{N})\geq 0$, the unmodified force $F(R)$ might be used, or other strategies employed to ensure movement towards a saddle point. A standard optimization algorithm (often L-BFGS in many implementations) then takes a step using this effective force to update the midpoint:
$R_{k+1}=\text{OptimizerStep}(R_{k},F_{trans}(R_{k}))$ (5)
The rotation and translation steps are repeated iteratively until convergence criteria, typically based on the magnitude of the total force $F(R)$, are met.
2.2 Statistical models
Quantitative assessment of saddle point search algorithm performance requires rigorous statistical methods, particularly when comparing multiple algorithms across diverse chemical systems. While foundational studies have provided crucial insights [2, 15, 21, 20], analyses often rely on qualitative comparisons or the eyeball norm and summary statistics based on relatively small datasets. Such approaches may not fully capture performance variability or allow for robust statistical inference regarding the significance and magnitude of performance differences. When larger system sets are used they are often hampered by point-estimates of error like the root mean square deviation [12, 37]. It is important to note that much of the saddle search literature for molecular systems makes several approximations valid for small angles or well behaved portions of the energy landscape, which may not hold in practice.
The number of potential energy surface (PES) calls (computational cost) represents count data, total computation time is typically positive and right-skewed, and search success is a binary outcome. These data types often violate the assumptions inherent in standard linear models (e.g., normality, homoscedasticity) [33]. Additionally, comparing algorithm performance typically involves applying multiple methods to the same set of chemical systems, resulting in repeated measures or clustered data. Observations from the same system are not independent, and failing to account for this structure can lead to inaccurate standard errors and potentially flawed conclusions. Variability arising from different software implementations or environments can also present challenges [22], underscoring the need for reproducible analyses.
To analyze the saddle search results, we focus on interpreting three algorithmic concerns, the computational cost, the total time taken, and the success ratio. These are considered in the context of varying the optimizer used in the rotation phase (CG versus LBFS) and the effect of using rotation removal (yes versus no) thus forming 4 sets of measurements on the 500 systems. We employed Bayesian generalized linear mixed-effects models [25]. This framework is well-suited to the data structure and research questions. GLMMs explicitly model the appropriate response distribution for each outcome (counts, continuous time, binary success) using link functions, while incorporating random effects to account for the repeated measures structure (i.e., baseline differences between systems). The Bayesian approach provides full posterior distributions for parameter estimates, facilitating comprehensive uncertainty quantification via credible intervals and enabling principled model comparison [4]. We use brms [4] to construct a Stan model [6] which is compiled into C++ code for analysis with the no U turn sampler (NUTS) [17]. These models are so efficient that they are run on a local machine with CPU configuration of 13th Gen Intel i7-1365U (12) @ 5.200GHz with 30 GB of available memory. Details of model fitting, prior specification, and convergence diagnostics are provided in the Supplementary Information.
2.2.1 Effect structures considered
All statistical models employed shared a common approach to constructing the linear predictor, denoted by $\eta_{ij}$, which corresponds to observation $i$ on system $j$. The expected value of the outcome ($\mu_{ij}$) is related to this linear predictor via an appropriate link function $g(\cdot)$, i.e., $g(\mu_{ij})=\eta_{ij}$, with the specific link function and response distribution detailed below for each outcome type.
Each linear predictor $\eta_{ij}$ included a random intercepts $u_{j}$, to account for the repeated measures within system $j$, i.e. $u_{j}\sim\text{Normal}(0,\sigma^{2}_{u})$.
The fixed-effects component was varied systematically to address different analytical questions. The specific fixed-effects structures examined were:
$\text{Fixed Effects}_{ij}=\begin{cases}\beta_{0}+\beta_{1}\text{DR}_{i(j))}&\text{(RotOptimizer)}\\ \beta_{0}+\beta_{2}\text{RR}_{i(j))}&\text{(RotRemoval)}\\ \beta_{0}+\beta_{1}\text{DR}_{i(j))}+\beta_{2}\text{RR}_{i(j))}+\beta_{3}(\text{DR}_{i(j))}\times\text{RR}_{i(j))})&\text{(Full)}\end{cases}$ (6)
Here, $\beta_{k}$ represents fixed-effects coefficients (intercept and slopes), $\text{DR}_{i(j))}$ is an indicator variable for the dimer rotation optimizer used (CG or L-BFGS) for observation $i$ within system $j$, and $\text{RR}_{i(j))}$ is an indicator variable for rotation removal (’no’ or ’yes’) within system $j$. The $\text{DR}_{i(j))}\times\text{RR}_{i(j))}$ term represents their interaction.
The first two effect settings focus on the main effect of a single factor (RotOptimizer or RotRemoval, respectively). When fitted to the complete dataset, these estimate the effect of the included factor while marginalizing over the levels of the excluded factor. Alternatively, as detailed in the Results for specific targeted comparisons, these structures might be applied to subsets of the data (e.g., comparing optimizers only when rotatons are removed).
The third effect setting considers the main effects of the dimer rotation optimizer and the rotation removal. The fourth effect setting represents the full interaction model. It simultaneously estimates the main effects of both factors and their interaction term. This model generally utilizes data from all four conditions per system (where available) to provide the most comprehensive assessment of how the factors jointly influence the outcome.
This approach allows for both focused investigation of individual factor effects, potentially mirroring simpler experimental designs, and a detailed analysis of their interplay using the full statistical power of the dataset within the multilevel framework. The specific response distribution and link function for each outcome metric are detailed in the following sections.
2.2.2 Modeling algorithmic efficiency
The primary figure of merit for saddle search algorithms, particularly when introducing new methodologies, is the number of calls made to the underlying potential energy surface (PES). This metric, referred throughout this manuscript as “PES calls,” directly correlates with computational cost, especially when high-level electronic structure methods are employed.
PES calls represent count data [33], which typically violate the assumptions of standard linear models, which assume the response variable is normally distributed. Beyond this, the difficulty of applying frequentist measures of statistical significance like the p-value for applied situations is fraught with controversy [26, 23]. As demonstrated in the Supplementary Information, even linear mixed-effects models, which account for per-molecule variability, can produce misleading results when applied directly to count data without appropriate transformations or distributional assumptions.
For comparing PES calls, we employed Bayesian generalized linear mixed-effects models (GLMMs) with a negative binomial response distribution with a logarithmic link function, which has support on the non negative integers. This distribution is appropriate for overdispersed count data like PES calls [33]. We thus use:
[…]
$\log(\mu_{ij})$ $=\text{Fixed Effects}_{ij}+u_{j}$
where $\text{PESCalls}_{ij}$ represents the observed PES calls for the $i^{th}$ observation on system $j$; $\mu_{ij}$ is the expected value; $\phi$ is the shape parameter of the negative binomial distribution; Fixed effects Eq. 6 comprise of the overall intercept and effect of the method; and $u_{j}$ is a random intercept for chemical system $j$, accounting for the paired nature of the data and modeled as $u_{j}\sim\text{Normal}(0,\sigma^{2}_{u})$.
2.2.3 Modeling total time
Total time analysis for saddle search algorithms is often perilous, since the bulk of compute intesive calculation are offloaded to external codes like VASP, ORCA etc. which may have strong scaling and depend heavily on the specific functional form of the energy function along with the runtime environment. Additionally, compared to the cost of making PES calls, the time spent within the saddle search algorithm is almost always negligible. Nevertheless, since total wall time elapsed is of practical importance, it provides a valid algorithmic constraint for decision making.
For estimating the effect on total computation time, a similar GLMM structure was employed. Given that time is a continuous, positive, and often right-skewed variable, we used a Gamma response distribution with a logarithmic link function:
$\text{TotalTime}_{ij}$ $\sim\text{Gamma}(\mu_{ij},\alpha)$ (8)
$\log(\mu_{ij})$ $=\text{Fixed Effects}_{ij}+u_{j}$
where $\text{TotalTime}_{ij}$ is the computation time for observation $i$ on system $j$, $\mu_{ij}$ is the expected time, $\alpha$ is the shape parameter of the Gamma distribution (inversely related to the variance when the variance is presented in terms of the squared mean). The linear predictor $\log(\mu_{ij})$ incorporates the fixed effects specified in Eq. 6 and the system-specific random intercept $u_{j}$, which is modeled as $u_{j}\sim\text{Normal}(0,\sigma^{2}_{u})$.
2.2.4 Modeling success probabilities
To analyze the factors influencing the likelihood of a successful saddle point search convergence, we fitted a Bayesian generalized linear mixed model appropriate for binary outcomes. This involved using a Bernoulli response distribution and a logit link function:
$\text{Success}_{ij}$ $\sim\text{Bernoulli}(p_{ij})$ (9)
$\text{logit}(p_{ij})$ $=\log\left(\frac{p_{ij}}{1-p_{ij}}\right)=\text{Fixed Effects}_{ij}+u_{j}$
Here, $\text{Success}_{ij}$ is a binary indicator for observation $i$ on system $j$, $p_{ij}$ is the probability of success, $\text{logit}(p_{ij})$ is the log-odds of success, and the linear predictor incorporates the fixed effects specified in Eq. (6) and the system-specific random intercept $u_{j}$, which is modeled as $u_{j}\sim\text{Normal}(0,\sigma^{2}_{u})$.
This model allows us to estimate how the different factors influence the odds of a successful search, while accounting for baseline differences in success rates across systems.
3 Applications
To demonstrate the utility of leveraging a Bayesian approach to performance statistices, we analyse the performance of the dimer method variants on the dataset of 500 initial configurations for small gas-phase organic molecules (7-25 atoms) introduced by hermesSellaOpenSourceAutomationFriendly2022 [15]. Saddle point searches were conducted using the EON software package [9]  [^1] Using the Github variant: https://github.com/TheochemUI/eOn, interfaced with NWChem [1] as the quantum mechanical engine. Energies and forces were calculated at the Hartree-Fock (HF) level of theory using RHF (restricted HF) for singlets and UHF (unrestricted HF) for doublets [19] with the 3-21G basis set. Saddle searches algorithms were considered converged when the maximum component of the atomic forces fell below 0.01 eV/Å. Comprehensive details regarding computational parameters, including self-consistent field thresholds, specific EON settings, and interface configurations, are provided in the Supplementary Information.
All models involved pairwise comparisons of dimer control modalities, using data filtered to include only systems where both methods in the pair successfully converged, and with the same saddle, with an energy difference of less than 0.01eV. Full interaction models are also fit. Since both PES calls and total time are proxies for computational effort, only the PES calls are presented here, with the total time model in the Supplementary.
3.1 Data overview
We applied the four dimer method variants (combinations of CG/L-BFGS rotation optimizers and enabling/disabling rotation removal) to the benchmark set of 500 initial configurations. An initial exploratory analysis of the performance metrics provides context for the subsequent Bayesian modeling, focusing here on the number of potential energy surface (PES) calls for successful runs and overall convergence success.
Figure 1: Binary convergence outcomes for dimer method calculations across 500 systems, ordered by median root-mean-squared-displacement (RMSD) of the intial geometry relative to the saddle point geometry. Each panel directly compares two settings, with the vertical position and color indicating the outcome. Gray points indicate both compared settings succeed. Yellow points indicate where both compared settings fail. (a) Comparing LBFGS (blue) and CG (red) with rotation removal enabled, both methods (gray) succeeded in 94.0% of systems, while CG additionally succeeded (red) in 3.6% (LBFGS in 0.6%, blue) and both failed (yellow) in 1.8%. (b) Without rotation removal, comparing LBFGS (blue) to CG (red) yielded 90.8% joint success (gray), with 5.8% (red) unique CG successes (LBFGS in 1.6%, blue) while the joint failure is 4.2% (yellow). (c) Contrasting rotation removal (blue) with no rotation removal (red) using the CG-rotations shows high joint success (95.8%, gray) with 0.8% additional (red) without removal (1.8% with rotation removal, blue) and 1.6% cases where both fail (yellow). (d) For the LBFGS optimizer, comparing across rotation removal (blue) resulted in 90.6% joint success (gray), without rotation removal (red) 1.8% additional successes while with rotation removal (blue) 4% additional cases succeed, and there are 3.6% joint failures (yellow). Failures, particularly the unique failures with LBFGS without rotation removal, are concentrated at higher RMSD values.
Figures 2 and 3 illustrate the difference in PES calls required for convergence, comparing pairs of calculations that reached the same saddle point (within 0.01 eV). When rotation removal was disabled (Fig. 2, 440 comparable systems), the CG-rotations dimer required fewer PES calls than L-BFGS-rotations in a majority of cases (55%, median reduction 21 calls), although L-BFGS is rather comparable in performance (in 43% cases, median reduction 16 calls). With CG-rotations (Fig. 3, 465 comparable systems), disabling rotation removal was markedly more efficient, requiring fewer PES calls than enabling it in over 99% of systems (median reduction 126 calls). Both figures indicate that performance differences tend to be larger for systems starting with higher initial RMSD values relative to the saddle point.
Figure 2: (a) Difference in the number of potential energy surface (PES) calls between L-BFGS and conjugate gradient (CG) rotations for dimer calculations that converged to the same saddle, defined as having a final energy difference of less than 0.01 eV. The x-axis represents the system index, sorted by the median RMSD. The y-axis displays the difference in PES calls, for changing the optimizer for the rotation of the dimer when rotation removal is not used. Consequently, negative values indicate that the CG algorithm required fewer PES calls (i.e., CG performed better), while positive values indicate that L-BFGS required fewer PES calls (i.e., LBFGS performed better). The median improvement is 21 calls occuring in 55.23% systems for CG-rotations compared to a median improvement of 16 calls in 43.18% systems for LBFGS-rotations across 440 systems. Systems which are farther away from the saddle point show a larger improvement with CG-rotations. (b) For the systems shown in (a), their log(median RMSD) plotted against System Index, with Median RMSD displayed on the y-axis.
Figure 3: (a) Difference in the number of potential energy surface (PES) calls between using rotation removal (T) and not using rotation removal (F) for dimer calculations that converged to the same saddle, defined as having a final energy difference of less than 0.01 eV. The x-axis represents the system index, sorted by the median RMSD. The y-axis displays the difference in PES calls, using CG-rotations with and without rotation removal. Consequently, negative values indicate that the rotation removal required fewer PES calls (i.e., it performed better), while positive values indicate turning off the rotation removal took fewer PES calls (i.e., no rotation removal performed better). The median improvement is 126 calls occuring in 99.35% systems for rotation removal not being used compared to a median improvement of 22 calls in 0.65% systems when rotation removal is used across 465 systems. Systems which are farther away from the saddle point show a larger improvement without rotation removal. (b) For the systems shown in (a), their log(median RMSD) plotted against System Index, with Median RMSD displayed on the y-axis.
Beyond the computational cost for successful runs, the overall convergence success is critical. Figure 1 presents a comparison of binary success outcomes across all 500 systems for which both methods in each comparison pair were run. High rates of joint success (where both compared methods converged) were generally observed across conditions. However, differences favouring the CG optimizer’s robustness are apparent. When comparing optimizers directly, CG achieved unique success (succeeding when LBFGS failed) considerably more often than LBFGS did, both with rotation removal enabled (3.6% vs 0.6%, panel a) and disabled (5.8% vs 1.6%, panel b). Disabling rotation removal slightly reduced joint success rates and increased joint failures from 1.8% to 4.2% when comparing the optimizers. Examining the impact of rotation removal itself highlights CG’s robustness further: using the CG optimizer (panel c), success rates remained very high (~96% joint success) with minimal sensitivity, as disabling rotation removal resulted in unique success in only 0.8% of cases. Conversely, using the LBFGS optimizer (panel d), disabling rotation removal significantly impacted performance; while joint success was still high (90.6%), unique successes for the enabled rotation removal setting were more frequent (4.0% vs 1.8% for disabled), and the joint failure rate rose notably to 3.6%. This suggests CG is less perturbed by the absence of rotation removal. As observed previously, convergence failures tend to concentrate at higher initial RMSD values, particularly in scenarios involving LBFGS without rotation removal.
This initial overview highlights apparent performance trends and differences between the method variants. However, these direct comparisons and visual inspections do not fully quantify the uncertainty associated with these effects or rigorously account for system-specific variability and potential interactions between the optimizer choice and rotation removal. Therefore, we employ the Bayesian hierarchical models detailed in Section 2.2 for a more robust statistical analysis of these metrics.
3.2 Effort analysis model
To quantify the effects of the dimer rotation optimizer and rotation removal on computational effort, we analyzed the number of PES calls using the Bayesian negative binomial generalized linear mixed-effects model described in Section 2.2 (Eq. 7). We fitted models corresponding to the different fixed-effects structures defined in Eq. 6, examining main effects within relevant data subsets as well as the full interaction model using all comparable data.
[…]
$\log(\mu_{ij})$ $=\eta_{ij}$
$\eta_{ij}$ =$=\begin{cases}\beta_{0}+\beta_{1}\text{DR}_{i(j))}+u_{j}&\text{(a)}\\ \beta_{0}+\beta_{2}\text{RR}_{i(j))}+u_{j}&\text{(b)}\\ \beta_{0}+\beta_{1}\text{DR}_{i(j))}+\beta_{2}\text{RR}_{i(j))}+\beta_{3}(\text{DR}_{i(j))}\times\text{RR}_{i(j))})+u_{j}&\text{(c)}\end{cases}$
$u_{j}$ $\sim\text{Normal}(0,\sigma^{2}_{u})$
First, applying the RotOptimizer model structure (Eq. (10)a), specifically to calculations where rotation removal was disabled, we examin the effect of the rotation optimizer ($\beta_{1}$). Switching from the CG reference to the L-BFGS optimizer resulted in a small but statistically credible increase in the median number of PES calls by 2.6% (95% CrI: [1.1%, 4.1%]). This suggests a slight efficiency advantage, in terms of PES calls, for the CG optimizer under these conditions.
Next, using the RotRemoval model structure (Eq. (10)b) applied to calculations using the CG optimizer, we examine the effect of enabling rotation removal ($\beta_{2}$). Enabling rotation removal led to a substantial and credible increase in the median PES calls by 41.5% (95% CrI: [38.6%, 44.4%]) compared to disabling it. This indicates a significant computational cost penalty associated with enabling rotation removal when using the CG optimizer in this setup.
Finally, the full interaction model (Eq. (10)c), incorporating both factors and their interaction simultaneously across 1805 valid observations, corroborated the main effects observed in the subset models (posterior distributions shown in Figure 4). The estimated median increase in PES calls for L-BFGS relative to CG (when rotation removal is off, $\beta_{1}$) was 2.6% (95% CrI: [0.7%, 4.5%]), corresponding to a multiplicative factor credibly just above 1 (median 1.03). Similarly, the median increase for enabling rotation removal relative to disabling it (when using CG, $\beta_{2}$) was substantial at 44.2% (95% CrI: [41.6%, 46.8%]), clearly visualized by the distribution centered around a multiplicative factor of 1.44 in Figure 4. The 95% credible interval for the interaction term ($\beta_{3}$), visually overlapping 1 in Figure 4 (multiplicative factor median 1.01, 95% CrI [0.99, 1.04]), provides no strong evidence that the effect of the optimizer choice significantly depends on the rotation removal setting, or vice versa. Furthermore, this full model estimated the standard deviation of the random intercepts for each system to be substantial (median $\sigma_{u}$ = 0.63, 95% CrI: [0.59, 0.67] on the log scale), indicating significant system-specific variation in the baseline number of PES calls required for convergence, independent of the method variations.
These findings from the hierarchical models, which account for system variability via random intercepts ($\sigma_{u}\approx$ 0.63), reinforce the trends observed in the exploratory data analysis (Fig. 2 and Fig. 1). They quantify the consistent, albeit small, PES call advantage of CG over L-BFGS ($\beta_{1}$) and the considerable computational overhead incurred by enabling this particular rotation removal implementation ($\beta_{2}$) for this dataset. The model diagnostics, including posterior predictive checks and residual analyses (shown for individual models in the preceding code sections, further details in SI), indicate reasonable model fit for the purpose of comparing these relative effects, despite some remaining heteroscedasticity common in complex chemical datasets.
Figure 4: Posterior distributions of fixed effects on the multiplicative scale for PES calls from the full interaction model. The x-axis represents the factor by which expected PES calls change relative to the baseline (CG optimizer, rotation removal disabled), with the dashed line at x=1 indicating no effect. Distributions show the main effect of using L-BFGS (vs CG, when rotation removal is disabled), the main effect of enabling rotation removal (vs disabling, when using the CG optimizer), and their interaction. Thick lines represent 95% credible intervals with median dots. Enabling rotation removal shows a large multiplicative increase (~1.44), L-BFGS shows a small increase (~1.03), and the interaction credibly includes 1.
3.3 Success rate model
We next analyzed the factors influencing the probability of successful convergence using the Bayesian Bernoulli generalized linear mixed-effects model specified in Section 2.2 (Eq. 11). Effects are reported as Odds Ratios (OR), representing the multiplicative change in the odds of success compared to a reference condition. An OR ¿ 1 indicates increased odds, OR ¡ 1 indicates decreased odds, and a 95% Credible Interval (CrI) that excludes 1 suggests a statistically credible effect.
$\text{Success}_{ij}$ $\sim\text{Bernoulli}(p_{ij})$ (11)
$\text{logit}(p_{ij})$ $=\log\left(\frac{p_{ij}}{1-p_{ij}}\right)=\eta_{ij}$
$\eta_{ij}$ =$=\begin{cases}\beta_{0}+\beta_{1}\text{DR}_{i}+u_{j}&\text{(a)}\\ \beta_{0}+\beta_{2}\text{RR}_{i}+u_{j}&\text{(b)}\\ \beta_{0}+\beta_{1}\text{DR}_{i}+\beta_{2}\text{RR}_{i}+\beta_{3}(\text{DR}_{i}\times\text{RR}_{i})+u_{j}&\text{(c)}\end{cases}$
$u_{j}$ $\sim\text{Normal}(0,\sigma^{2}_{u})$
First, applying the RotOptimizer model structure (Eq. (11)a) specifically to the subset of calculations where rotation removal was disabled, we examined the effect of the optimizer ($\beta_{1}$). L-BFGS showed significantly lower odds of success compared to the CG reference. The estimated median OR for L-BFGS relative to CG was 0.3 (95% CrI: [0.13, 0.60]), indicating that the odds of convergence with L-BFGS were roughly only 30% of the odds with CG under these conditions. This model also revealed substantial variability between systems in their baseline odds of success (posterior median $s_{u}$ = 2.6 on the log-odds scale, 95% CrI: [1.67, 3.81]).
Next, using the RotRemoval model structure (Eq. (11)b) specifically to the subset of calculations using CG-rotations, we examined the effect of rotation removal ($\beta_{2}$). The median OR for enabling rotation removal versus disabling it was 2.0 (95% CrI: [0.73, 6.35]). Although the point estimate suggests doubled odds of success when enabling rotation removal, the wide credible interval comfortably includes 1. Consequently, this model provides no strong statistical evidence that enabling rotation removal significantly alters the odds of success compared to disabling it when using the CG optimizer. Substantial between-system variability in baseline success odds was again noted (median sd(Intercept) = 4.2, 95% CrI: [2.63, 6.91] from this model’s summary; the full model estimate is used below for consistency).
Finally, the full interaction model (Eq. (11)c), analyzing all 2000 observations across the four conditions, allowed for a detailed assessment of success odds. While the predicted probability of success for an average system is very high (¿0.98) under all conditions, as visualized in the main panel of Figure 5, the model revealed statistically credible differences when analyzing the underlying odds ratios as seen in the inset of Figure 5. It reinforced the substantially lower odds of success for L-BFGS compared to CG (median OR = 0.2, 95% CrI: [0.09, 0.45]); this relative difference, indicating higher robustness for CG, is subtly reflected in Figure 5 by the slight shift of the L-BFGS probability distributions (right column) towards lower values compared to CG (left column). The model also showed no credible main effect for rotation removal (median OR = 1.9, 95% CrI: [0.74, 5.07]) nor a credible interaction (median Interaction OR = 1.1, 95% CrI: [0.33, 3.42]), consistent with the high degree of similarity between the top (’no’) and bottom (’yes’) panels within each optimizer in Figure 5. Thus, while both optimizers perform well on average, the primary factor credibly influencing the relative odds of success in this analysis is the choice of optimizer, with CG demonstrating significantly higher odds of convergence than L-BFGS across both rotation removal settings in this dataset. The large estimated standard deviation for the system-specific random intercepts from the full model (posterior median $\sigma_{u}$ = 3.6, 95% CrI: [2.79, 4.80]) underscores that system-specific properties remain a major determinant of convergence success or failure for individual cases, explaining deviations from the high average success probability. Model diagnostics indicated good convergence and fit (details in SI).
Figure 5: Posterior distributions for convergence success derived from the full interaction model (brmspfidall). The main panel displays the predicted probability of success for an average system under the four method conditions (Optimizer-Rotation Removal, see legend), illustrating high probabilities (¿0.98) for all settings. The inset shows the posterior distributions for the fixed effects on the Odds Ratio (OR) scale relative to the baseline (CG, rotation removal disabled), with the dashed line at OR=1 indicating no change in odds. While absolute probabilities are high, the inset quantifies the relative effects, showing the OR for L-BFGS vs CG (bottom distribution) is credibly below 1, while the ORs for rotation removal (top) and the interaction (middle) credibly include 1. The distributions for CG are slightly shifted towards higher probabilities compared to L-BFGS. Distributions for enabling (’yes’) versus disabling (’no’) rotation removal are nearly identical within each optimizer column, but do tend to cause a rightward shift. Thick lines in the inset represent 95% credible intervals with median dots.
The preceding analyses provide a quantitative, statistically robust assessment of how the choice of dimer rotation optimizer and the application of rotation removal influence both the computational effort (PES calls) and the likelihood of success for saddle point searches within this benchmark set. Having established these effects and accounted for system-specific variability using Bayesian hierarchical models, we now turn to discussing the implications of these findings in the context of theoretical expectations, practical algorithm choice, and the broader methodology of benchmarking computational chemistry methods.
4 Discussion
Our analysis, employing Bayesian hierarchical models to account for system-specific variability, provides quantitative insights into the performance of CG and L-BFGS optimizers for dimer rotation, and the effect of enabling rotation removal, across a large benchmark set. The key findings indicate a small but credible PES call advantage and significantly higher convergence odds for the CG optimizer compared to L-BFGS. Conversely, enabling rotation removal incurred a substantial PES call penalty without providing a statistically credible improvement in success probability in the full model applied to this dataset.
The superior performance of CG over L-BFGS, particularly its enhanced robustness demonstrated by the success probability analysis (OR for L-BFGS vs CG $\approx$ 0.2 [0.09, 0.45]), aligns with potential theoretical advantages of CG in navigating complex potential energy surfaces. While modern L-BFGS methods incorporate strategies to ensure positive-definite Hessian approximations [32], they may still be susceptible to challenges in regions with complex curvature compared to methods like CG that directly seek movement along specific eigenmodes [24]. The observed difference in computational effort per successful run was modest (L-BFGS requiring ~2.6% [0.7%, 4.5%] more PES calls than CG when rotation removal was off), suggesting that the primary practical advantage of CG in this context lies in its higher probability of reaching convergence.
Counter-intuitively, enabling rotation removal, theoretically designed to simplify the optimization by projecting out external coordinates [27], resulted in a significant increase in computational cost (median +44% PES calls vs disabled) without a corresponding statistically credible improvement in success probability in our full model (OR for RotRem Yes vs No $\approx$ 1.9 [0.74, 5.07]). This discrepancy likely highlights the sensitivity of optimization performance to implementation details and the specific nature of the PES. The projection involved in rotation removal interacts with the optimizer’s steps and Hessian updates; for this dataset containing potentially fragmenting or aggregating systems near saddle points, where the separation of internal and external coordinates can be ill-defined [31], this projection may hinder convergence pathways that are accessible without it. Thus we observe less efficient trajectories when rotation removal is active for the majority of these systems.
This study also underscores the value of applying rigorous statistical modeling to computational benchmarks, moving beyond simple averages or qualitative comparisons. Traditional approaches can be easily confounded by the large system-to-system variability inherent in chemical datasets, as evidenced by the substantial standard deviation estimated for the random intercepts in our models i.e. $\sigma_{u}\approx$ 0.63 for PES calls (on the logarithmic scale), \$$\sigma$$\approx$\$3.6 for success odds(on the logit scale). Bayesian hierarchical models explicitly account for this variability through random effects ($u_{j}$). Furthermore, they leverage partial pooling (shrinkage), effectively borrowing strength across systems to produce more stable and reliable estimates for the effects of different methods, mitigating the influence of outlier systems - an effect sometimes colloquially related to “regression towards the mean”. This framework also provides direct quantification of uncertainty via full posterior distributions and credible intervals (e.g., Figs. 4, 5), facilitating more nuanced conclusions than point estimates or p-values alone [25, 26]. The ability to model interactions formally, although no strong interaction was detected here, is another advantage for dissecting performance factors.
There are very few results to imply why LBFGS for the rotation-phase would be of much use. The L-BFGS does not in most cases enjoy the same quadratic guarantee [3] as the BFGS, and additionally, the CG is less likely to be trapped by higher eigenpairs [24]. These results are not surprising in that respect, although it is clear that the difference is primarily in success probabilities, with no appreciable change in terms of computational performance.
Given the results, particularly the higher robustness of CG and the general inefficiency introduced by enabling rotation removal in this context, a practical strategy for high-throughput workflows might involve defaulting to the CG optimizer without rotation removal. A fallback mechanism could then activate rotation removal only for specific systems that initially fail to converge, potentially capturing benefits for problematic cases where external modes interfere, although identifying such cases a priori remains challenging. This “chain of methods” approach warrants further study but illustrates how statistically grounded benchmarks can inform practical workflow design beyond simply selecting a single “best” method.
5 Conclusions
We introduced and applied a Bayesian hierarchical modeling framework to rigorously analyze performance metrics from computational chemistry benchmarks, demonstrating its utility for comparing dimer method variants across a large, diverse dataset. This statistically robust approach moves beyond simple averages or visualizations, accounting for system-specific variability and providing nuanced uncertainty quantification, offering a more actionable view for benchmarking algorithms.
From the generalized linear mixed models (Section 2.2), using the LBFGS for rotations showed a small but statistically discernible increase of 3.4% (95% CrI: [1.1%, 5.8%]) in PES calls. Furthermore, the Bernoulli model suggests a substantial improvement in success probability for Dimer CG, with an estimated 23% higher success rate (95% CI: 5%, 40%), supporting the conclusion that CG-rotations are more robust.
We assert that these results, along with the automated capabilities of most workflow engines suggest a “chain” of methods approach rather than an absolute best. Although given the slight increase in performance and the significant increase in robustness it is hard to argue for the LBFGS rotations, the situation for the chemically intuitive external degree of freedom removal is more nuanced. We were able to comprehensively study the implementation of the rotation removal in EON. While the concept of removing external degrees of freedom is well-founded, our results indicate that, for the systems and implementation studied, this removal often increases the number of PES calls and reduces the likelihood of successful convergence, contrary to its intended purpose. There are however, situations where the rotation removal is required for convergence. Since the data demonstrates that for these random saddle point initializations that the rotation removal for such fragmenting systems is not optimal in most cases, it is more natural to first run CG-rotation without external rotation removal, and have a fallback rule to run CG-rotation with external rotation only for systems which otherwise fail. This strategy leverages the higher success rate of CG in the majority of cases while providing a fallback option for more challenging systems.
Our benchmark results and analysis are packaged in a user-friendly repository, showing integration with Snakemake for high throughput calculations and EON for the energy and forces calculation through NWChem along with R code for the figures and running the Bayesian analysis. It is also, to the best of our knowledge, the largest benchmark study done on the modalities of the dimer method. Our analysis provides a more nuanced and statistically robust assessment of performance than traditional approaches, which often rely on simple averages or visualizations that can obscure important details. We expect this reproducible resource will be indispensible to others in the community performing high throughput chemistry or algorithm development.
6 Supplementary Material
Model validations for the Bayesian models are fairly standard and verbose, so they are present in the accompanying supplementary material. The supplementary material also includes videos of the path taken to the saddle (or failure) for key systems referred to in the main text.
Acknowledgments
This work was supported by the Icelandic Research Fund (grant no. 217436-053).
RG thanks Dr. Miha Gunde, Dr. Amrita Goswami, Dr. Moritz Sallermann, Prof. Debabrata Goswami, Prof. Laurent Karim Béland, Prof. Morris Riedel, Mrs. Sonaly Goswami and Mrs. Ruhila Goswami for discussions. RG also thanks Prof. Normand Mousseau and Prof. Birgir Hrafnkelsson for manuscript suggestions and edits.
Conflict of interest
The authors declare no conflict of interest.